Four times Pete's age minus three times Jame's age equals 20.
Pete's age is also two times Jame's age.
How old is Pete?

Answer by fractalier(6550)

Call their ages, p and j.
Then we have
4p - 3j = 20
and
p = 2j
Now substitute this into the first equation and we get
4(2j) - 3j = 20
8j - 3j = 20
5j = 20
j = 4
p = 8